Calculate the energy (in MeV) released when ? decay converts uranium 232U (atomic mass = 232.037146 u) into thorium 228Th (atomic mass = 228.028731 u). The atomic mass of an ? particle is 4.002603 u.
Answer #1

Here ,


mass lost = 232.037146 - (228.028731 + 4.002603)

mass lost = 0.005812 amu

Now,

mass lost = 9.65105221
Answer #2

SOLUTION :


mass lost in decay 

= mu - mt - m particle 

= 232.037146 - 228.02873 - 4.002603 

= 0.005813 amu .

= 0.005813 * (1.6 * 10^(-27) )  kg         (1 amu = 1.6*10^(-27) kg) 

= 9.3008 * 10^(-30) kg 


Now, energy released , E 

= m c^2 

= 9.3008 * 10^(-30) * (3*10^8)^2        (c = speed of light = 3*10^8 m/s)

= 8.37072 * 10^(-13) kg.m/s^2 * m (= N-m = J)

= 8.3702 * 10^(-13) J

= 8.3702 * 10^(-13) / (1.6 * 10^(-19))  ev     ( 1 ev = 1.6*10^(-19) J)

= 5.231375 * 10^6 

= 5.231375 Mev (ANSWER).

answered by: Tulsiram Garg